OpenStudy (anonymous):

Solve for n: 2 3 = n 18

OpenStudy (tyteen4a03):

I cannot understand the question. Can you elaborate?

OpenStudy (anonymous):

is it 2/3=n/18?

OpenStudy (anonymous):

if so then n=2*18/3=2*6=12

OpenStudy (anonymous):

thanks man